# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ Conclusion
$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{=}(X)\}}{p(D\in\mathcal{S}_{=}(X))},&\ \text{if }X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{<}(X)\}}{p(D\in\mathcal{S}_{<}(X))},&\ \text{if }X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\\ Y_{\text{min}},&\ \text{if }X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{*}(X)\}}{p(D\in\mathcal{S}_{*}(X))},&\ \text{if }X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\\ Y_{\text{min}},&\ \text{if }X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}.\\ \end{cases}$
Thus, $L(\pi_{e})=\mathbb{E}[\psi_{L}(X,D,Y)]$.
@@
The proof for $U(\pi_{e})$ follows similarly. We directly write the following through linearity of expectations and three applications of Lemma B.1.
$U(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{=}(x)]$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{\not\geq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{*}(x)]$
$\quad+\mathbf{1}\{x\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{max}]$
$=\mathbb{E}[Y\frac{\mathbf{1}\{D\in\mathcal{S}_{=}(x)\}}{P(D\in\mathcal{S}_{=}(x))}\mathbf{1}\{X\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y\frac{\mathbf{1}\{D\in\mathcal{S}_{>}(x)\}}{P(D\in S_{>}(x))}\mathbf{1}\{X\in\mathcal{X}_{>}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y_{max}\mathbf{1}\{X\in\mathcal{X}_{\not\geq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y\frac{\mathbf{1}\{D\in S_{*}(x)\}}{P(D\in S_{*}(x))}\mathbf{1}\{X\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}$
$\quad+Y_{max}\mathbf{1}\{X\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}]$
@@
Because this different enumeration of the sets $(X\in\mathcal{X}_{=}\cap\pi_{e}\neq a_{0})$, $(X\in\mathcal{X}_{>}\cap\pi_{e}\neq a_{0})$, $(X\in\mathcal{X}_{\not\geq}\cap\pi_{e}\neq a_{0})$, $(X\in\mathcal{X}_{*}\cap\pi_{e}=a_{0})$, and $(X\not\in\mathcal{X}_{*}\cap\pi_{e}=a_{0})$ are also disjoint, only one product of the indicator functions inside the expectation above ever evaluates to $1$. Thus, we can equivalently express the expression inside the expectation above as the piecewise function
$\psi_{U}(X,D,Y)$
@@
$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{=}(X)\}}{p(D\in\mathcal{S}_{=}(X))},&\ \text{if }X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{>}(X)\}}{p(D\in\mathcal{S}_{>}(X))},&\ \text{if }X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\\ Y_{\text{max}},&\ \text{if }X\in\mathcal{X}_{\not\geq},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{*}(X)\}}{p(D\in\mathcal{S}_{*}(X))},&\ \text{if }X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\\ Y_{\text{max}},&\ \text{if }X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\\ \end{cases}$
Thus, $U(\pi_{e})=\mathbb{E}[\psi_{U}(X,D,Y)]$.

### v2

@@ Conclusion
$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\}}{P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X))},&\ \text{if }\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\\ Y_{\text{min}},&\ \text{if }\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{\Pi\in\mathbf{\Pi}^{e}(X)\}}{P(\Pi\in\mathbf{\Pi}^{e}(X))},&\ \text{if }\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\\ Y_{\text{min}},&\ \text{if }\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\end{cases}$
Thus, $L(\pi_{e})=\mathbb{E}[\psi_{L}(Y,X,\Pi)]$.
@@
The proof for $U(\pi_{e})$ follows similarly. We directly write the following through linearity of expectations and two applications of Lemma D.1.
$U(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)]$
$\quad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(X)]$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{max}]$
$=\mathbb{E}[Y\frac{\mathbf{1}\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\}}{P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X))}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y_{max}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+\frac{\mathbf{1}\{\Pi\in\mathbf{\Pi}^{e}(X)\}}{P(\Pi\in\mathbf{\Pi}^{e}(X))}\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}$
$\quad+Y_{max}\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}]$
@@
Because this different enumeration of the sets $(\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\pi_{e}\neq a_{0})$, $(\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\pi_{e}\neq a_{0})$, $(\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}=a_{0})$, and $(\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}=a_{0})$ is also disjoint, only one product of the indicator functions inside the expectation above ever evaluates to $1$. Thus, we can equivalently express the expression inside the expectation above as the piecewise function
$\psi_{U}(Y,X,\Pi)$
@@
$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\}}{P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X))},&\ \text{if }\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\\ Y_{\text{max}},&\ \text{if }\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{\Pi\in\mathbf{\Pi}^{e}(X)\}}{P(\Pi\in\mathbf{\Pi}^{e}(X))},&\ \text{if }\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\\ Y_{\text{max}},&\ \text{if }\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\end{cases}$
Thus, $U(\pi_{e})=\mathbb{E}[\psi_{U}(Y,X,\Pi)]$.